OpenStudy (anonymous):

as the shuttle bus comes to a sudden stop to avoid hitting a dog it accelerates uni formally at -4.1ms/s^2 as it slows from 9.0 ms/s to 0.0 m/s find the time interval for acceleration for this bus

OpenStudy (anonymous):

2.2 v = vi +a (delta)t 0= 9 + -4.1t -9= 4.1t -9/ -4.1 = t 2.2 =t

OpenStudy (anonymous):

thank you so much i just got this right now still a bit confused though